OpenStudy (anonymous):

The area of the parallelogram is 32 sq miles. Find its base and height. (There is this parallelogram and the height it is x-2 and the width is x+2

jigglypuff314 (jigglypuff314):

area of parallelogram = (base)(height) plug in what you know 32 = (x-2)(x+2) foil, simplify, and factor

OpenStudy (anonymous):

So would the answer be x=6,-6?

jigglypuff314 (jigglypuff314):

yes x would equal that, but what was the question asking for?

OpenStudy (anonymous):

It was asking the base and height.

jigglypuff314 (jigglypuff314):

base = x + 2 and height = x - 2 so now that you have x, plug in... but keep in mind that distances/lengths have to be positive